# AI Agent Behavioral Science

### v1

@@ Abstract
Recent advances in large language models (LLMs) have enabled AI systems to behave in increasingly human-like ways, exhibiting planning, adaptation, and social dynamics across increasingly diverse, interactive, and open-ended scenarios. These behaviors are not solely the product of the models’ internal architecture, but emerge from their integration into agentic systems that operate within situated contexts, where goals, feedback, and interactions shape behavior over time. This shift calls for a new scientific lens: AI Agent Behavioral Science. Rather than focusing only on internal mechanisms, this paradigm emphasizes the systematic observation of behavior, design of interventions to test hypotheses, and theory-guided interpretation of how AI agents act, adapt, and interact over time. We systematize a growing body of research across individual, multi-agent, and human-agent interaction settings, and further demonstrate how this perspective informs responsible AI by treating fairness, safety, interpretability, accountability, and privacy as behavioral properties. By unifying recent findings and laying out future directions, we position AI Agent Behavioral Science as a necessary complement to traditional approaches, providing essential tools for understanding, evaluating, and governing the real-world behavior of increasingly autonomous AI systems.
@@ 1 Introduction
Recent advances in large language models (LLMs) have profoundly transformed how we build and interact with AI systems (See Figure 1). Beyond static prediction tasks, LLMs are now embedded into interactive systems that simulate agents capable of reasoning wei2022emergent , planning huang2022language , and adaptation yao2023react . For instance, when placed in a virtual village, LLM agents develop routines, hold conversations, and even organize a Valentine’s Day party park2023generative . In social deduction games like Werewolf or Avalon, they engage in deception, persuasion, and alliance formation xu2023exploring ; lan2023llm . These behaviors are not pre-programmed, but emerge through situated interaction, and evolve in response to other agents, human users, and feedback from the environment. As such deployments proliferate, they open up a timely opportunity: to study AI systems not merely as statistical models, but as behavioral entities whose actions, adaptations, and social patterns can be empirically observed and systematically understood.
@@
Traditional approaches to understanding AI have focused on internal mechanisms: architectures nguyen2020wide , weights erhan2010does , attention patterns clark2019does , and training objectives radford2019language (see Table 1). These model-centric views, inspired by fields like physics and neuroscience, have yielded deep insights into what models encode and how they process information. However, they rest on the assumption that behavior can be determined from within. While this may hold in static and well-bounded tasks, it breaks down in socially embedded and open-ended environments krishna2022socially , where behavior is shaped not just by internal computation, but by interaction history, social context, and feedback loops. Model-centric tools can surface latent capacities, but they struggle to explain the emergence of complex behaviors such as negotiation, coordination, and deception. Crucially, such behaviors rarely emerge from the AI model alone. Rather, they arise when AI models are embedded in agentic systems, i.e., architectures that incorporate memory, planning, tool use, and action modules weng2023llm , transforming static models into dynamic, interactive entities. In this light, the model is to behavior what the brain is to action: a substrate that enables but does not determine. Just as human behavior cannot be understood in isolation from environment and experience clark1998being ; anderson2003embodied , AI behavior must be studied as a product of not only system design but also situated interaction.
@@
We frame this emerging perspective as the AI Agent Behavioral Science paradigm, i.e., the study of how AI agents act, adapt, and interact in situated contexts. Drawing inspiration from human and animal behavioral research, this paradigm emphasizes systematic observation of behavior, hypothesis-driven intervention design, and theory-informed interpretation to uncover agent behavioral regularities and mechanisms. It asks not only what models can do in principle, but what agents actually do in practice, and more specifically, how behavioral patterns emerge, stabilize, generalize, or misalign over time given specific roles, incentives, environments, and peers. While much current research focuses on LLM-based agents, the core questions generalize to any AI system capable of goal-directed interaction, whether symbolic, embodied, or multimodal. Importantly, this paradigm also unlocks new pathways for advancing responsible AI dignum2019responsible , as fairness, safety, interpretability, accountability, and privacy are not just properties of models, but consequences of behavior—of how systems act under pressure, adapt over time, and interact with others in the wild.
@@
This paradigm builds upon several foundational works. Rahwan et al. rahwan2019machine call for a science of machine behavior that treats AI systems as empirical subjects of behavioral study. Mei et al. mei2024turing demonstrate how behavioral science tools can be repurposed to assess LLM preferences and traits, drawing comparisons with a global dataset of human behavior. The emerging term “AI Behavioral Science” has been used in both scholarly comments meng2024ai and dedicated venues lakkaraju2024first to reflect this paradigm shift. While these works outline the promise of a behavioral approach, they are largely conceptual or programmatic. By contrast, our survey takes a step further by organizing this perspective into a coherent research paradigm, systematizing emerging empirical findings, and identifying shared methods, dimensions, and open questions. We also situate this work within broader sociotechnical conversations about AI in society. Tsvetkova et al. tsvetkova2024new propose a new sociology of human-machine systems, viewing hybrid networks of people and AI agents as complex systems with emergent dynamics. Brinkmann et al. brinkmann2023machine explore machine culture, emphasizing how AI systems increasingly participate in generating and transmitting cultural patterns. These views reinforce the idea that AI systems should not only be engineered and interpreted, but also observed, evaluated, and governed as participants in social ecosystems.
@@
In this survey, we aim to lay the groundwork for a scientific understanding of AI behavior. Our contribution is twofold: first, we conceptualize AI Agent Behavioral Science as a coherent research paradigm—one that complements model-centric analysis by shifting the focus toward interaction, adaptation, and emergent dynamics of AI agents. Second, we synthesize a growing body of work on LLM agents to highlight how behavioral patterns can be observed, measured, and theorized across different contexts. For the rest of the paper, Section 2, 3, and 4 examine three key domains of behavior: individual agent dynamics, multi-agent interactions, and human-agent interactions. Section 5 reviews how AI agent behavior can be adapted, re-interpreted and organized with the Fogg Behavior Model. Section 6 applies the behavioral lens to responsible AI, highlighting how ethical principles are behaviorally measured and optimized. Finally, Section 7 outlines critical open questions and charts promising future directions for this emerging field.
@@
2 Emergent Individual AI Agent Behaviors
In the era of large language models, understanding the behavior of LLM-agent systems, particularly their decision-making processes, has become a central focus of research. Based on the social cognitive theory bandura2001social , the characteristics of AI agent behavior can be systematically analyzed through three key dimensions: intrinsic attributes, environmental constraints, and behavioral feedback (Figure 2). These dimensions collectively provide a comprehensive framework for exploring how decision-making emerges and evolves in intelligent agents: (1) Intrinsic attributes: An agent’s behavior is shaped by internal traits such as emotions, cognitive patterns, value judgments, and biases. These intrinsic factors determine how an agent processes information and makes decisions, forming the foundation of its behavioral tendencies. (2) Environmental constraints: The environment imposes constraints that shape an agent’s behavior. Cultural norms, geographical factors, and institutional rules define boundaries and influence value judgments. Agents must navigate these constraints to align their actions with societal expectations and regulations. (3) Behavioral feedback: Decision-making is significantly shaped by social interactions, where agents adjust their behaviors in response to external feedback and observed outcomes in dynamic environments. This continuous process of adaptation is driven by the influence of others’ actions and the resulting consequences on individual or group behavior. Following this framework, a detailed exploration of how decision-making behavior emerges in LLM-based agents is presented below (see Table 2 for summary).
@@ 2 Emergent Individual AI Agent Behaviors
Table 2: Summary of emergent individual AI agent behaviors.
@@ 2.1 Intrinsic Attributes: Intrinsic Traits and Decision Mechanisms
The research on the fundamental characteristics of AI agents can be divided into three main areas: (1) emotions and cognition: exploring how LLMs simulate emotional responses and cognitive processes, which are essential for enhancing their human-like interactions. (2) economic rationality: investigating how LLMs make decisions that mimic rational behavior in decision theory and game theory contexts. (3) bias: examining how LLMs may inadvertently reflect societal biases and the potential consequences of such biases on fairness and decision-making in AI applications.
@@ 2.3 Behavioral Feedback: Social Influence and Relationship Construction
This primarily refers to self-play, with AlphaGo Silver2017MasteringTG being a highly representative study. The research explores whether AI can autonomously learn the game of Go solely through self-interaction and feedback, without relying on any human knowledge. Ultimately, after extensive self-play, AlphaGo is able to surpass human decision-making in gameplay, defeating world champions in Go competitions. Moreover, it is capable of developing strategies that human players have never used before.
@@
In multi-agent competitive and cooperative scenarios, feedback from other agents influences the behavior of a single agent. Agents in interactive environments may actively cooperate or seek confrontation. Agents can spontaneously form cooperation through dynamic multi-agent interactions, even without explicit instructions in competitive scenarios Wu2024ShallWT . In cooperative relationships, agents allocate goals and avoid conflicts, while in competitive relationships, agents take deceptive actions against opponents and engage in active confrontation 10.5555/3295222.3295385 . Dai et al. dai2024artificial establish a multi-agent sandbox simulation where agents initially adopt zero-sum competitive behaviors. As agents interact and receive feedback from one another, they gradually learn to cooperate and form a social contract.
@@
This section primarily includes two aspects: exploring or guiding AI agent behavior through human feedback at each step and during strategic interactions. In classical behavioral economics games, AI’s decision-making is influenced by human observation and demographic factors. When an agent is asked to explain the choices or told that its choices will be observed by a third party, it becomes significantly more generous. Moreover, when the agent knows the human player’s gender, it tends to be more selfish in its allocations. And AI agents also exhibit significant changes in behaviors as they experience different roles in a game mei2024turing .
AI demonstrates greater rationality in complex strategic interactions with humans by relying more on modeling and optimization. In investment interactions, AI compensates disadvantaged players based on their relative contributions and penalizes free riders, achieving a favorable balance between productivity (surplus) and equality (Gini coefficient) Koster2022HumancenteredMD . In the game of diplomacy, an AI agent can predict other players’ responses and adjust its strategy accordingly. It does not blindly trust other players’ proposals; instead, it makes decisions based on its own interests Bakhtin2022HumanlevelPI .
@@ 2.4 Summary
In this section, we review the decision-making behavior of single agent through the lens of social cognitive theory, focusing on three key dimensions: intrinsic attributes, environmental constraints, and behavioral feedback. Intrinsic attributes reveal that LLMs exhibit human-like emotions, cognitive patterns, and biases, with models like GPT-4 demonstrating advanced capabilities in emotional recognition, implicit reasoning, and theory of mind, though challenges remain in achieving consistent economic rationality. Environmental constraints highlight the importance of cultural adaptation and institutional alignment, showing that LLMs perform better when tailored to culturally specific contexts and legal frameworks, but still face issues such as social identity bias and sensitivity to contradictory information. Behavioral feedback underscores the dynamic adaptation mechanisms of agents in interactive scenarios, including self-interaction, multi-agent cooperation or competition, and human-AI interactions, where AI demonstrates strategic rationality, cooperative tendencies, and adaptability to feedback. Collectively, these insights provide a comprehensive understanding of how decision-making emerges and evolves in intelligent agents, emphasizing both their potential and limitations in complex environments.
@@
Looking ahead, several critical directions need further exploration. First, many current evaluations are limited in scale and scenario diversity, which constrains the generalizability of findings. Developing richer and more representative benchmarks is essential to ensure the validity and robustness of results across different contexts. Second, addressing the ”black-box” nature of large language models remains a pressing challenge. Achieving greater transparency and controllability in model behavior will enhance the interpretability and generalization of outcomes, paving the way for more reliable applications. Overall, research into the individual behavior of LLM agents is still in its early stages but holds immense promise. With continued advancements, this field has the potential to unlock groundbreaking insights and applications, offering a fertile ground for future exploration.
3 Emergent Multi-agent Behaviors
When multiple individuals interact, new and complex behaviors can emerge that go beyond the capabilities or intentions of any single individual. As AI systems become increasingly autonomous and socially embedded, understanding the dynamics of multi-agent behaviors becomes essential not only for ensuring safety and alignment, but also for unlocking the potential of large-scale coordination, creativity, and decision-making. We organize this section into three primary behavioral patterns observed in LLM-driven multi-agent systems, as shown in Figure 3. The first concerns cooperative dynamics, where agents align toward shared goals through mechanisms such as deliberation, role coordination, and norm-following. The second explores competitive dynamics, in which agents pursue conflicting objectives, leading to behaviors such as deception, retaliation, or resource hoarding. The third examines open-ended interaction dynamics, where agents are free to define their own goals and social relationships, resulting in emergent structures such as institutions, shared routines, or consensus cultures. Table 3 summarizes the important literature along these three patterns, and notes the emergent behavior observed in each study.
@@ 3.4 Summary (Chen Lin)
4 Emergent AI Agent Behaviors in Human-Agent Interaction
As AI agents become increasingly embedded in human-centered environments, their interactions with humans give rise to distinct behavioral patterns radivojevic2024llms . These behaviors are not merely outcomes of model architecture or training objectives, but are shaped by the roles agents come to occupy in the situated social environments tsvetkova2024new ; krishna2022socially . Some of these roles are explicitly assigned—for instance, an AI assistant may be designed to exhibit self-disclosure to foster trust traeger2020vulnerable . Others emerge through dynamic interaction, as agents adapt to human preferences, social signals, or adversarial pressures. Regardless of origin, roles structure the way AI agents behave in relation to humans: how they communicate, influence, co-create, or contest. In this section, we examine the kinds of behavior that emerge when AI agents inhabit particular roles in human-AI interactions. We group these roles into two broad contexts: In cooperative contexts, AI agents support aligned human goals by adapting to social cues, stimulating exploration, or reshaping group structures. In rivalrous contexts, AI agents engage in competition or exert asymmetric influence, pursuing objectives that may conflict with those of human users. We summarize relevant research in Table 4.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have enabled the development of AI agents that exhibit increasingly human-like behaviors, including planning, adaptation, and social dynamics across diverse, interactive, and open-ended scenarios. These behaviors are not solely the product of the internal architectures of the underlying models, but emerge from their integration into agentic systems operating within specific contexts, where environmental factors, social cues, and interaction feedbacks shape behavior over time. This evolution necessitates a new scientific perspective: AI Agent Behavioral Science. Rather than focusing only on internal mechanisms, this perspective emphasizes the systematic observation of behavior, design of interventions to test hypotheses, and theory-guided interpretation of how AI agents act, adapt, and interact over time. We systematize a growing body of research across individual agent, multi-agent, and human-agent interaction settings, and further demonstrate how this perspective informs responsible AI by treating fairness, safety, interpretability, accountability, and privacy as behavioral properties. By unifying recent findings and laying out future directions, we position AI Agent Behavioral Science as a necessary complement to traditional model-centric approaches, providing essential tools for understanding, evaluating, and governing the real-world behavior of increasingly autonomous AI systems.
@@ 1 Introduction
Recent advances in large language models (LLMs) have profoundly transformed how we build and interact with AI systems, particularly through the emergence of AI agents (See Figure 1). An AI agent is an autonomous system that perceives its environment and takes actions to achieve certain goals xi2025rise . For instance, when placed in a virtual village, LLM-based AI agents develop routines, hold conversations, and even organize a Valentine’s Day party park2023generative . In social deduction games like Werewolf or Avalon, they engage in deception, persuasion, and alliance formation xu2023exploring ; lan2023llm . These behaviors are not pre-programmed, but emerge through situated interaction, and evolve in response to other agents, human users, and feedback from the environment. As such deployments proliferate, they open up a timely opportunity: to study AI systems not merely as statistical models, but as behavioral entities whose actions, adaptations, and social patterns can be empirically observed and systematically understood in context.
@@
Traditional approaches to understanding AI have focused on internal mechanisms: architectures nguyen2020wide , weights erhan2010does , attention patterns clark2019does , and training objectives radford2019language (see Table 1). These model-centric views, inspired by fields like physics and neuroscience, have yielded deep insights into what models encode and how they process information. However, they rest on the assumption that behavior can be determined and fully understood from within. However, as AI models grow increasingly complex, pinpointing which specific components or neurons trigger particular behaviors has become challenging. Moreover, in socially embedded and open-ended environments krishna2022socially , behavior is shaped not just by internal computation, but by interaction history, social context, and feedback loops. Model-centric tools struggle to explain the emergence of complex behaviors such as negotiation schneider2024negotiating , coordination zhang2023exploring , and deception xu2023exploring . Crucially, such behaviors rarely emerge from the AI model alone. Rather, they arise when AI models are embedded in agentic systems, i.e., architectures that incorporate memory, planning, tool use, and action modules weng2023llm , transforming static models into dynamic, interactive entities. In this light, the model is to behavior what the brain is to action: a substrate that enables but does not determine. Just as human behavior cannot be understood in isolation from environment and experience clark1998being ; anderson2003embodied , AI agent behavior must be studied as a product of not only system design but also situated interaction.
@@
We frame this emerging perspective as the AI Agent Behavioral Science paradigm, i.e., the study of how AI agents act, adapt, and interact in situated contexts. Drawing inspiration from human and animal behavioral research, this paradigm emphasizes systematic observation of behavior, hypothesis-driven intervention design, and theory-informed interpretation to uncover agent behavioral regularities and mechanisms. It asks not only what models can do in principle, but what agents actually do in practice, and more specifically, how behavioral patterns emerge, stabilize, generalize, or misalign over time given specific roles, incentives, environments, and peers. While much current research focuses on LLM-based agents, the core questions generalize to any AI system capable of goal-directed interaction, whether symbolic, embodied, or multimodal. Importantly, this paradigm also unlocks new pathways for advancing responsible AI dignum2019responsible , reframing fairness, safety, interpretability, accountability, and privacy from static and one-shot properties of models to dynamic and context-dependent attributes.
@@
This paradigm builds upon several foundational works. Rahwan et al. rahwan2019machine call for a science of machine behavior that treats AI systems as empirical subjects of behavioral study. Mei et al. mei2024turing demonstrate how behavioral science tools can be repurposed to assess LLM preferences and traits, drawing comparisons with a global dataset of human behavior. Both scholarly comments meng2024ai and dedicated venues lakkaraju2024first have started to recognize the importance of behavioral science toward understanding and building AI agents. While these works outline the promise of a behavioral approach, they are largely conceptual or programmatic. By contrast, our paper takes a step further by organizing this perspective into a coherent research paradigm, systematizing emerging empirical findings, and identifying shared methods, dimensions, and open questions. We also situate this work within broader sociotechnical conversations about AI in society. Tsvetkova et al. tsvetkova2024new propose a new sociology of human-machine systems, viewing hybrid networks of people and AI agents as complex systems with emergent dynamics. Brinkmann et al. brinkmann2023machine explore machine culture, emphasizing how AI systems increasingly participate in generating and transmitting cultural patterns. These views reinforce the idea that AI systems should not only be engineered and interpreted, but also observed, evaluated, and governed as participants in social ecosystems.
In this paper, we aim to lay the groundwork for a scientific understanding of AI agent behavior. Our contributions are summarized below:
  • We conceptualize AI Agent Behavioral Science as a coherent research paradigm—one that complements model-centric analysis by shifting the focus toward interaction, adaptation, and emergent dynamics of AI agents.
  • We synthesize a growing body of work on AI agents to highlight how behavioral patterns can be observed, measured, theorized, and adapted across individual agent, multi-agent, and human-agent interaction scenarios.
  • We discuss how AI Agent Behavioral Science offers new possibilities to achieve responsible AI, for both measurement and optimization.
  • We propose six promising research directions in this area.
@@
For the rest of the paper, Section 2, 3, and 4 examine AI agent behavior in individual agent, multi-agent, and human-agent interaction settings, respectively. Section 5 focuses on the adaptation and optimization of AI agent behavior, interpreting and understanding existing methods under the Fogg Behavior Model fogg2009behavior . Section 6 applies the behavioral lens to responsible AI, highlighting how ethical principles are behaviorally measured and optimized. Finally, Section 7 outlines critical open questions and promising research directions for this emerging field.
@@
2 Behavioral Foundations of Individual AI Agents
Just as human individuals constitute the basic units of human societies, the behaviors of individual AI agents form the foundational layer for modeling higher-order interaction patterns and collective dynamics. Therefore, we begin by examining the behavioral foundations at the level of individual AI agents. Drawing inspiration from the social cognitive theory (SCT) bandura2001social , we organize existing research around three key dimensions that shape agent behavior over time: intrinsic attributes, environmental constraints, and behavioral feedback (Figure 2). These dimensions offer a structured foundation for analyzing how situated decision-making and adaptation emerge in AI agents:
  • Intrinsic attributes shape agent behavior through internal traits such as emotions, cognitive patterns, value judgments, and biases, which determine how agents process information and make decisions.
  • Environmental constraints influence agent behavior through external structures such as cultural norms, geographical contexts, and institutional rules, which define boundaries and social expectations.
  • Behavioral feedback captures how agents continuously adapt their actions in response to social interaction, external feedback, and observed consequences.
Following this framework, we now introduce a range of emergent behaviors at the individual agent level (see Table 2 for summary).
@@ 2 Behavioral Foundations of Individual AI Agents
Table 2: Summary of emergent individual agent behaviors.
@@ 2.1 Intrinsic Attributes: Intrinsic Traits and Decision Mechanisms
The research on the intrinsic attributes of AI agents can be divided into three main areas: (1) emotions and cognition: exploring how LLMs simulate emotional responses and cognitive processes, which are essential for enhancing their human-like interactions. (2) economic rationality: investigating how LLMs make decisions that mimic rational behavior in decision theory and game theory contexts. (3) bias: examining how LLMs may inadvertently reflect societal biases and the potential consequences of such biases on fairness and decision-making in AI applications.
@@ 2.3 Behavioral Feedback: Social Influence and Relationship Construction
This primarily refers to self-play, with AlphaGo Silver2017MasteringTG being a highly representative study. The research explores whether an AI agent can autonomously learn the game of Go solely through self-interaction and feedback, without relying on any human knowledge. Ultimately, after extensive self-play, AlphaGo is able to surpass human decision-making in gameplay, defeating world champions in Go competitions. Moreover, it is capable of developing strategies that human players have never used before.
@@
In multi-agent scenarios, feedback from other agents influences the behavior of a single agent. Agents in interactive environments may actively cooperate or seek confrontation. In cooperative contexts, agents allocate goals and avoid conflicts, while in competitive contexts, agents take deceptive actions against opponents and engage in active confrontation 10.5555/3295222.3295385 . Moreover, agents can spontaneously form cooperation through dynamic multi-agent interactions, even without explicit instructions in competitive scenarios Wu2024ShallWT . Dai et al. dai2024artificial establish a multi-agent sandbox simulation where agents initially adopt zero-sum competitive behaviors. As agents interact and receive feedback from one another, they gradually learn to cooperate and form a social contract.
@@
This section primarily includes two aspects: exploring or guiding AI agent behavior through human feedback at each step and during strategic interactions. In classical behavioral economics games, agents’ decision-making is influenced by human observation and demographic factors. When an agent is asked to explain the choices or told that its choices will be observed by a third party, it becomes significantly more generous. Moreover, when the agent knows the human player’s gender, it tends to be more selfish in its allocations. And AI agents also exhibit significant changes in behaviors as they experience different roles in a game mei2024turing .
AI agents demonstrate greater rationality in complex strategic interactions with humans by relying more on modeling and optimization. In investment interactions, an AI agent compensates disadvantaged players based on their relative contributions and penalizes free riders, achieving a favorable balance between productivity (surplus) and equality (Gini coefficient) Koster2022HumancenteredMD . In the game of diplomacy, an AI agent can predict other players’ responses and adjust its strategy accordingly. It does not blindly trust other players’ proposals; instead, it makes decisions based on its own interests Bakhtin2022HumanlevelPI .
@@ 2.4 Summary
In this section, we examine the behavioral foundations of individual AI agents through three SCT-inspired lenses: intrinsic attributes, environmental constraints, and behavioral feedback. Focusing on LLM-powered AI agents, we find that they demonstrate striking human-like capabilities in cognitive reasoning, emotion recognition, and theory of mind, though they still fall short in consistent economic rationality and remain sensitive to task framing. Environmentally, these agents show partial cultural adaptability and context awareness, with improved performance when aligned to local knowledge and institutional norms, but they remain prone to biases and contradictions in politically or ethically sensitive scenarios. Under behavioral feedback, LLM-powered agents have shown dynamic adaptation in self-play, agent-agent interaction, and agent-human interaction, adopting cooperative, strategic, or even manipulative behaviors in response to feedback and social context. Collectively, these findings offer a grounded picture of how behavior emerges from the interplay between internal mechanisms and external conditions.
@@
Nevertheless, many current evaluations are limited in scale and scenario diversity, which constrains the generalizability of findings. Developing richer and more representative benchmarks is essential to ensure the validity and robustness of results across different contexts. Besides, addressing the ”black-box” nature of large language models remains a pressing challenge. Achieving greater transparency and controllability in model behavior will enhance the interpretability and generalization of outcomes, paving the way for more reliable applications.
3 Behavioral Dynamics in Multi-Agent Interactions
When multiple individuals interact, new and complex behaviors can emerge that go beyond the capabilities or intentions of any single individual. Having examined the behavioral foundations of individual AI agents, we now turn to the dynamics that emerge when multiple such agents interact. In this section, we conceptually hold agents’ intrinsic traits constant and focus on how social interaction and environmental structure give rise to higher-order behaviors. We organize these behaviors into three primary patterns, differentiated by the goal relationships among agents (see Figure 3).
  • Cooperative dynamics emerge when agents pursue shared or aligned goals, often facilitated by deliberation, role coordination, and norm-following;
  • Competitive dynamics arise when agents have conflicting goals, leading to behaviors such as deception, retaliation, or strategic exclusion;
  • Open-ended interaction dynamics occur when agents act with independent, evolving, or non-specific goals, allowing for the spontaneous emergence of institutions, routines, and social structures.
Table 3 summarizes the important literature along these three patterns, and notes the emergent behavior observed in each study.
@@ 3.4 Summary
4 Behavioral Roles of AI Agents in Human Interactions
As AI agents become increasingly embedded in human-centered environments, their interactions with humans give rise to distinct behavioral patterns radivojevic2024llms . These behaviors are not merely outcomes of model architecture or training objectives, but are shaped by the roles agents come to occupy in the situated social environments tsvetkova2024new ; krishna2022socially . Some of these roles are explicitly assigned, e.g., an AI assistant may be designed to exhibit self-disclosure to foster trust traeger2020vulnerable ; Others emerge through dynamic interaction, as agents adapt to human preferences, social signals, or adversarial pressures. Regardless of origin, roles structure the way AI agents behave in relation to humans: how they communicate, influence, co-create, or contest. In this section, we examine the kinds of behavior that emerge from interactions with humans when AI agents inhabit particular roles. We group these roles into two broad contexts:
  • In cooperative contexts, AI agents support aligned human goals by adapting to social cues, stimulating exploration, or reshaping group structures.
  • In rivalrous contexts, AI agents engage in competition or exert asymmetric influence, pursuing objectives that may conflict with those of human users.
We summarize representative studies in Table 4, and will detail them below.